OpenStudy (anonymous):

Write an equation of the line that passes through the points (-3,4) and (-5,1).

OpenStudy (stefrheart):

You should first find the slope and that would help alot y2-y1/x2-x1

OpenStudy (stefrheart):

So what would the slope be?

OpenStudy (anonymous):

-6?

OpenStudy (stefrheart):

okay so \[\frac{ 1-4 }{ -5--3 }\]

OpenStudy (anonymous):

-3/-8

OpenStudy (anonymous):

-3/-2

OpenStudy (stefrheart):

yes so that becomes 3/2 because the negatives cancel each other out

OpenStudy (stefrheart):

Next you need to do point slope form y1-y=m(x-x1)

OpenStudy (stefrheart):

do you understand?

OpenStudy (triciaal):

@stefrheart be careful with the slope formula. the direction is important the change in y / CORRESPONDING change in x y1 -y = m( x1 - x)

OpenStudy (stefrheart):

No I am correct @triciaal on the formula that the y goes first

OpenStudy (triciaal):

@Stefrheart you clearly misunderstood what I wrote!!! even look at what you posted which is correct and the same as mine. the coordinates are (x, y) and (x1, y1) same direction/ corresponding y - y1 x - x1 using (y x ) (y1 x1) what you had was from above " Next you need to do point slope form y1-y=m(x-x1)" you have y1 - y x - x1 using (y1 x) and (y x1) can you see it now ? if not plot the points and review for a second.